OpenStudy (anonymous):

Solve this system: x+2y=5 2x+4y=2

OpenStudy (mathmale):

You have several different methods of solution to choose from: 1. elimination by addition or subtraction 2. elim. by substitution 3. graphing ... and several more.... Which would YOU like to try first?

OpenStudy (anonymous):

Erm,I've been trying elimination uvu

OpenStudy (mathmale):

So, you've been trying elimination by substitution, or elimination by addition/subtraction? Which?

OpenStudy (anonymous):

addition/subtraction

OpenStudy (mathmale):

OK. Let's look at our system: x+2y=5 2x+4y=2 I'd like for you to take the first equation and multiply every term in it by -2. Please type your result here.

OpenStudy (texaschic101):

he could just reduce the second equation

OpenStudy (mathmale):

True, and I will suggest he heed your advice, later. Thanks.

OpenStudy (texaschic101):

my bad...I will not butt in again...sorry

OpenStudy (anonymous):

-2x-4y=-20

OpenStudy (mathmale):

wouldn't that be -10 instead of -20? What do you see here?

OpenStudy (anonymous):

Sorry, I meant -10 but I'm trying to rush uAu

OpenStudy (mathmale):

OK x+2y=5 2x+4y=2 becomes -2x-4y = -10 2x+4y=2 Please combine all the like terms. In other words, combine -2x and 2x to obtain zero. what about the y terms? what about the constants on the right side of these equations?

OpenStudy (anonymous):

That was my problem, they both are eliminated! But if you plug it in, then it's as if it has no solution

OpenStudy (anonymous):

And -10 doesnt equal 2

OpenStudy (mathmale):

That's exactly right. Since 0 on the left side cannot equal 2-10 on the right, this system of linear equations has no solution. that's it. End of story.

OpenStudy (mathmale):

Note that we're not checking whether or not -10 equals 2; instead, we're combining those quantities; result is -8, right? 0 is not equal to -8, so our system has NO SOLUTION.

OpenStudy (anonymous):

Thank youso much! I thought it would be no solution but I was doubting myself. uvu

OpenStudy (mathmale):

I'm so glad we could work together and that this effort confirmed your initial solution. See you again on OpenStudy!